40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 5 of 326  Not logged in ELOG logo
ID Date Author Type Category Subject
  16206   Wed Jun 16 19:34:18 2021 KojiUpdateGeneralHVAC

I made a flow sensor with a stick and tissue paper to check the airflow.

- The HVAC indicator was not lit, but it was just the blub problem. The replacement bulb is inside the gray box.

- I went to the south arm. There are two big vent ducts for the outlets and intakes. Both are not flowing the air.
  The current temp at 7pm was ~30degC. Max and min were 31degC and 18degC.

- Then I went to the vertex and the east arm. The outlets and intakes are flowing.

Attachment 1: HVAC_Power.jpeg
HVAC_Power.jpeg
Attachment 2: South_Arm.jpeg
South_Arm.jpeg
Attachment 3: South_End_Tenperature.jpeg
South_End_Tenperature.jpeg
Attachment 4: Vertex.jpeg
Vertex.jpeg
Attachment 5: East_Arm.jpeg
East_Arm.jpeg
  16205   Wed Jun 16 17:24:29 2021 YehonathanUpdateCDSOpto-isolator for c1auxey

I updated the wiring diagram according to Koji's suggestion. According to the isolator manual, this configuration requires that the isolator input be configured as PNP.

Additionally, when the switch in the coil driver is open the LED in the isolator is signaling an on-state. Therefore, we might need to configure the Acromag to invert the input.

There are the Run/Aquire channels that we might need to add to the wiring diagram. If we do need to read them using slow channels, we will have to pull them up like the EnableMon channels to use them like in the wiring diagram.

Attachment 1: Optical_isolator_Wiring.pdf
Optical_isolator_Wiring.pdf
  16204   Wed Jun 16 13:20:19 2021 Anchal, PacoSummaryCamerasMon 7 in Control Room Replaced

We replaced the Mon 7 with an LCD monitor from back bench. It is fed the analog signal from BNC converted into VGS with a converter box that Paco bought. We can replace this monitor with another monitor if it is required on the back bench. For now, we definitely need a monitor to show IMC camera's up there.

Attachment 1: IMG_20210616_083810.jpg
IMG_20210616_083810.jpg
  16203   Tue Jun 15 21:48:55 2021 KojiUpdateCDSOpto-isolator for c1auxey

If my understanding is correct, the (photo receiving) NPN transistor of the optocoupler is energized through the acromag. The LED side should be driven by the coil driver circuit. It is properly done for the "enable mon" through 750Ohm and +V. However, "Run/Acquire" is a relay switch and there is no one to drive the line. I propose to add the pull-up network to the run/acquire outputs. This way all 8 outputs become identical and symmetric.

We should test the configuration if this works properly. This can be done with just a manual switch, R=750Ohm, and a +V supply  (+18V I guess).

Attachment 1: Acromag_RTS_BI_config.jpg
Acromag_RTS_BI_config.jpg
  16202   Tue Jun 15 15:26:43 2021 Anchal, PacoSummaryAUXXend Green Laser PDH OLTF measurement loop algebra, excitation at control point

Attachment 1 shows the case when excitation is sent at control point i.e. the PZT output. As before, free running laser noise \eta in units of Hz/rtHz is added after the actuator and I've also shown shot noise being added just before the detector.

Again, we have a access to three output points for measurement. \alpha right at the output of mixer (the PDH error signal), \beta the feedback signal to be applied by uPDH box (PZT Mon) and \gamma the output of the summing box SR560.

Doing loop algebra as before, we get:

\large \alpha = \frac{\eta}{K(s) A(s)} \frac{G_{OL}(s)}{1 - G_{OL}(s)} + \frac{\chi}{C(s) K(s) A(s)} \frac{G_{OL}(s)}{1 - G_{OL}(s)} - \frac{\nu_e}{K(s) } \frac{G_{OL}(s)}{1 - G_{OL}(s)}

\large \beta = \frac{\eta}{A(s)} \frac{G_{OL}(s)}{1 - G_{OL}(s)} + \frac{\chi}{C(s) A(s)} \frac{G_{OL}(s)}{1 - G_{OL}(s)} - \nu_e \frac{G_{OL}(s)}{1 - G_{OL}(s)}

\large \gamma= \frac{\eta}{A(s)} \frac{G_{OL}(s)}{1 - G_{OL}(s)} + \frac{\chi}{C(s) A(s)} \frac{G_{OL}(s)}{1 - G_{OL}(s)} - \nu_e \frac{1}{1 - G_{OL}(s)}

So measurement of \large G_{OL}(s) can be done by

\large G_{OL}(s) \approx \frac{\beta}{\gamma}

  • For frequencies, where \large G_{OL}(s) is large enough, to have an SNR of 100, we need that ratio of \large \nu_e to integrated noise is 100.
  • Assuming you are averaging for 'm' number of cycles in your swept sine measurement, time of integration for the noise signal would be \large \frac{m}{f}where f is the frequency point of the seeping sine wave.
    • This means, the amplitude of integrated laser frequency noise at either \large \beta or \large \gamma would be \large \sqrt{\left(\frac{\eta(f)}{A(f)}\right)^2\frac{f}{m}} = \frac{\eta(f) \sqrt{f}}{A(f)\sqrt{m}}
    • Therefore, signal to laser free running noise ratio at f would be \large S = \frac{\nu_eA(f)\sqrt{m}}{\eta(f) \sqrt{f}}.
    • This means to keep a constant SNR of S, we need to shape the excitation amplitude as \large \nu_e \sim S \frac{\eta(f) \sqrt{f}}{A(f)\sqrt{m}}
    • Putting in numbers for X end Green PDH loop, laser free-running frequency noise ASD is 1e4/f Hz/rtHz, laser PZT actuation is 1MHz/V, then for 10 integration cycles and SNR of 100, we get: \large \nu_e \sim 100 \times \frac{10^4 \sqrt{f}}{f \times10^6 \sqrt{10}} = \frac{30\, mV}{\sqrt{f}}
  • Assuming you are averaging for a constant time \large \tau in swept sine measurement, then the amplitude of integrated laser free noise would be \large \sqrt{\left(\frac{\eta(f)}{A(f)}\right)^2 \frac{1}{\tau}} = \frac{\eta(f) }{A(f)\sqrt{\tau}}
    • In this case, signal to laser free-running noise ratio at f would be \large S = \frac{\nu_eA(f)\sqrt{\tau}}{\eta(f)}
    • This means to keep a constant SNR of S, we need to shape the excitation amplitude as \large \nu_e \sim S\frac{\eta(f)}{A(f)\sqrt{\tau}}
    • Again putting in numbers as above and integration time of 1s, we need an excitation amplitude shape \large \nu_e \sim 100 \times \frac{10^4 }{f \times10^6 \sqrt{1}} = \frac{1\, V}{f}

This means at 100 Hz, with 10 integration cycles, we should have needed only 3 mV of excitation signal to get an SNR of 100. However, we have been unable to get good measurements with even 25 mV of excitation. We tried increasing the cycles, that did not work either.

This post is to summarize this analysis. We need more tests to get any conclusions.

Attachment 1: AuxPDHloop.pdf
AuxPDHloop.pdf
  16201   Tue Jun 15 11:46:40 2021 Ian MacMillanUpdateCDSSUS simPlant model

I have added more degrees of freedom. The model includes x, y, z, pitch, yaw, roll and is controlled by a matrix of transfer functions (See Attachment 2). I have added 5 control filters to individually control UL, UR, LL, LR, and side. Eventually, this should become a matrix too but for the moment this is fine.

Note the Unit delay blocks in the control in Attachment 1. The model will not compile without these blocks.

Attachment 1: x1sup_isolated-6-15-v1.pdf
x1sup_isolated-6-15-v1.pdf
Attachment 2: C1_SUS_SINGLE_PLANT-6-15-v1.pdf
C1_SUS_SINGLE_PLANT-6-15-v1.pdf
  16200   Mon Jun 14 18:57:49 2021 AnchalUpdateAUXXend is unbearably hot. Green laser is loosing lock in 10's of seconds

Working in Xend with mask on has become unbearable. It is very hot there and I would really like if we fix this issue.


Today, the Xend Green laser was just unable to hold lock for longer than 10's of seconds. The longest I could see it hold lock was for about 2 minutes. I couldn't find anything obviously wrong with it. Attached are noise spectrums of error and control points. The control point spectrum shows good matching with typical free running laser noise.

Are the few peaks above 10 kHz in error point spectrum worrysome? I need to think more about it in a cooler place to make sure.

I wanted to take a high frequency spectrum of error point to make sure that higher harmonics of 250 kHz modulation frequency are not leaking into the PDH box after demodulation. However, the lock could not be maintained long enough to take this final measurement. I'll try again tomorrow morning. It is generally cooler in the mornings.


This post is just an update on what's happening. I need to work more to get some meaningful inferences about this loop.

Attachment 1: XAUX_PDH_Err_In_ASD.pdf
XAUX_PDH_Err_In_ASD.pdf
Attachment 2: XAUX_PZT_Out_Mon_ASD.pdf
XAUX_PZT_Out_Mon_ASD.pdf
  16199   Mon Jun 14 15:31:30 2021 YehonathanUpdateCDSOpto-isolator for c1auxey

I checked the BI situation on the HAM-A coil driver. It seems like these are sinking BIs and indeed need to be isolated from the Acromag unit GND to avoid contamination.

The BIs will have to be isolated on a different isolator. Now, the wires coming from the field (red) are connected to the second isolator's input and the outputs are connected to the Acromag BI module and the Acromag's RTN.

I updated the wiring diagram (attached) and the wiring spreadsheet.

In the diagram, you can notice that the BI isolator (the right one) is powered by the Acromag's +15V and switched when the coil driver's GND is supplied. I am not sure if it makes sense or not. In this configuration, there is a path between the coil driver's GND and the Acromag's GND but its resistance is at least 10KOhm. The extra careful option is to power the isolator by the coil driver's +V but there is no +V on any of the connectors going out of the coil driver.

I installed an additional isolator on the DIN rail and wired the remaining BOs (C1:SUS-ETMY_SD_ENABLE, C1:SUS-ETMY_LR_ENABLE) through it to the DB9F-4 feedthrough. I also added DB9F-3 for incoming wires from the RTS and made the required connection from it to DB9F-4.

I tested the new isolated BOs using the Windows machine (after stopping Modbus). As before, I measure the resistance between pin 5 (coil driver GND) and the channel under test. When I turn on the BO I see the resistance drops from inf to 166ohm and back to inf when I turn it off. Both channels passed the test.

 

Attachment 1: Optical_isolator_Wiring.pdf
Optical_isolator_Wiring.pdf
  16198   Fri Jun 11 20:19:50 2021 KojiSummaryBHDBHD OMC invacuum wiring

Stephen and I discussed the in-vacuum OMC wiring.

- One of the OMCs has already been completed. (Blue)
- The other OMC is still being built. It means that these cables need to be built. (Pink)
- However, the cables for the former OMC should also be replaced because the cable harness needs to be replaced from the metal one to the PEEK one.
- The replacement of the harness can be done by releasing the Glenair Mighty Mouse connectors from the harness. (This probably requires a special tool)
- The link to the harness photo is here: https://photos.app.goo.gl/3XsUKaDePbxbmWdY7

- We want to combine the signals for the two OMCs into three DB25s. (Green)
- These cables are custom and need to be designed.

- The three standard aLIGO-style cables are going to be used. (Yellow)

- The cable stand here should be the aLIGO style.

Attachment 1: 40mBHD_OMC_wiring.pdf
40mBHD_OMC_wiring.pdf
  16197   Thu Jun 10 14:01:36 2021 AnchalSummaryAUXXend Green Laser PDH OLTF measurement loop algebra

Attachment 1 shows the closed loop of Xend Green laser Arm PDH lock loop. Free running laser noise gets injected at laser head after the PZT actuation as \eta. The PDH error signal at output of miser is fed to a gain 1 SR560 used as summing junction here. Used in 'A-B mode', the B port is used for sending in excitation \nu_e e^{st} where s = i\omega.

We have access to three ports for measurement, marked \alpha at output of mixer, \beta at output of SR560, and \gamma at PZT out monitor port in uPDH box. From loop algebra, we get following:

\large \left[ (\alpha - \nu_e) K(s)A(s) + \eta \right ]C(s)D(s) = \alpha

\large \Rightarrow (\alpha - \nu_e) G_{OL}(s) + \eta C(s)D(s) = \alpha, where \large G_{OL}(s) = C(s) D(s) K(s) A(s) is the open loop transfer function of the loop.

\large \Rightarrow \alpha = \eta \frac{C(s) D(s)}{1 - G_{OL}(s)} \quad -\quad \nu_e\frac{G_{OL}(s)}{1 - G_{OL}(s)}

\large \Rightarrow \beta = \eta \frac{C(s) D(s)}{1 - G_{OL}(s)} \quad -\quad \nu_e\frac{1}{1 - G_{OL}(s)}

\large \Rightarrow \gamma = \eta \frac{1}{K(s)} \frac{G_{OL}(s)}{1 - G_{OL}(s)} \quad -\quad \nu_e\frac{K(s)}{1 - G_{OL}(s)}

So measurement of \large G_{OL}(s) can be done in following two ways (not a complete set):

  1. \large G_{OL}(s) \approx \frac{\alpha}{\beta} = \frac{G_{OL}(s) - \frac{\eta C(s)D(s)}{\nu_e}}{1 - \frac{\eta C(s)D(s)}{\nu_e}}, if excitation amplitude is large enough such that \large \frac{\eta C(s)D(s)}{\nu_e} \ll 1over all frequencies.
    • In this method however, note that SR785 would be taking ratio of unsuppresed excitation at \large \alpha with suppressed excitation at \large \beta.
    • If the closed loop gain (suppression) \large 1/(1 - G_{OL}(s))is too much, the excitation signal might drop below noise floor of SR785 while measuring \large \beta.
    • This would then appear as a flat response in the transfer function.
    • This happened with us when we tried to measure this transfer function using this method. Below few hundered Hz, the measurement will become flat at around 40 dB.
    • Increasing the excitation amplitude where suppression is large should ideally work. We even tried to use Auto level reference option in SR785.
    • But the PDH loop gets unlocked as soon as we put exciation above 35 mV at this point in this loop.
  2. \large \frac{G_{OL}(s)}{K(s)} \approx \frac{\alpha}{\gamma} = \frac{G_{OL}(s) - \frac{\eta C(s)D(s)}{\nu_e}}{K(s)\left(1 - \frac{\eta C(s)D(s)}{\nu_e}\right )}, if excitation amplitude is large enough such that \large \frac{\eta C(s)D(s)}{\nu_e} \ll 1over all frequencies.
    • In this method, channel 1 (denominator) on SR785 would remain high in amplitude throughout the measurement avoiding the above issue of suppression below noise floor.
    • We can easily measure the feedback transfer funciton \large K(s) with the loop open. Then multiplying the two measurements should give us estimate of open loop transfer function.
    • This is waht we did in 16194. But we still could not increase the excitation amplitude beyond 35 mV at injection point and got a noisy measurement.
    • We checked yesterday coherence of excitation signal with the three measurment points \large \alpha, \beta, \gamma and it was 1 throughout the frequency region of measurement for excitation amplitudes above 20 mV.
    • So as of now, we are not sure why our signal to noise was so poor in lower frequency measurement.
Attachment 1: AUX_PDH_LOOP.pdf
AUX_PDH_LOOP.pdf
  16196   Wed Jun 9 18:29:13 2021 Anchal, PacoSummaryALSCheck for saturation in ITMX SOS Driver

We did a quick check to make sure there is no saturation in the C1:SUS-ITMX_LSC_EXC analog path. For this, we looked at the SOS driver output monitors from the 1X4 chassis near MC2 on a scope. We found that even with 600 x 10 = 6000 counts of our 619 Hz excitation these outputs in particular are not saturating (highest mon signal was UL coil with 5.2 Vpp). In comparison, the calibration trials we have done before had 600 counts of amplitude, so we can safely increase our oscillator strength by that much yes


Things that remain to be investigated -->

  • What is the actual saturation level?
  • Two-tone intermodulation?
  16195   Wed Jun 9 13:50:48 2021 Ian MacMillanUpdateCDSSUS simPlant model

I have attached an updated transfer function graph with the residual easier to see. I thought here I would include a better explanation of what this transfer function was measuring.

This transfer function was mainly about learning how to use DTT and Foton to make and measure transfer functions. Therefore it is just measuring across a single CDS filter block. X1SUP_ISOLATED_C1_SUS_SINGLE_PLANT_Plant_POS_Mod block to be specific. This measurement shows that the block is doing what I programmed it to do with Foton. The residual is probably just because the measured TF had fewer points than the calculated one.

The next step is to take a closed-loop TF of the system and the control module.

After that, I want to add more degrees of freedom to the model. both in the plant and in the controls.

Attachment 1: SingleSusPlantTF.pdf
SingleSusPlantTF.pdf
  16194   Wed Jun 9 11:46:01 2021 Anchal, PacoSummaryAUXXend Green Laser PDH OLTF measurement

We measured the Xend green laser PDH Open loop transfer function by following method:

  • We first measured the feedback transfer function 'K' directly.
    • See attachment 2 for this measurement. We measured Out2/exc here.
  • Then, we closed the loop as shown in attachment 1with SR560 as a summing juntion at error point.
    • We injected excitation through B channel in SR560 and measured transfer function Out1/Out2.
    • This measurement should give us G_{OL} / K by loop alegbra.
  • Then we multiplied the two transfer function measurements to get open loop transfer function.

Result:

  • Our measurement gives the same UGF of 10kHz and phase margin of 53.5 degrees as reported in 13238.
  • The shape of measurement also follows 1/f above 10 Hz atleast.
  • Our measurement might not be correct below 10 Hz but we did not see any saturation or loss of lock in 1Hz to 10 Hz measurement.
  • This OLTF is different from the modelled OLTF here even though the UGF matches.
  • The feedback gain is supposed to roll-off faster than 1/f in 30Hz to 1kHz region but it does not seem to in our measurement.
  • This suggests that the actual uPDH box is shaping the loop different from what schematic suggests. This might mean that the gain is much lower in the low frequency region than we would like it to be.
  • We will investigate the reason of difference between model and measurement unless someone has a better explaination for the descripancy.
Attachment 1: image-6f2923a3-01ce-4d04-bc53-d8db0238e195.jpg
image-6f2923a3-01ce-4d04-bc53-d8db0238e195.jpg
Attachment 2: image-72223f4b-3b74-4574-a7ad-de6628a2c5e9.jpg
image-72223f4b-3b74-4574-a7ad-de6628a2c5e9.jpg
Attachment 3: X_Green_ARM_PDH_OLTF.pdf
X_Green_ARM_PDH_OLTF.pdf
  16193   Tue Jun 8 11:54:39 2021 YehonathanUpdateCDSBI channels on c1auxey

I tested the digital inputs the following way: I connected a DB9 breakout to DB9M-5 and DB9M-6 where digital inputs are hosted. I shorted the channel under test to GND to turn it on.

I observed the channels turn from Disabled to Enabled using caget when I shorted the channel to GND and from Enabled to Disabled when I disconnected them.

I did this for all the digital inputs and they all passed the test.

I am still waiting for the other isolator to wire the rest of the digital outputs.

Next, I believe we should take some noise spectra of the Y end before we do the installation.

Quote:

Tomorrow I will test the BIs using EPICs.

 

  16192   Tue Jun 8 11:40:53 2021 Anchal, PacoSummaryALSSingle Arm Actuation Calibration with IR ALS Beat

We attempted to simulate "oscillator based realtime calibration noise monitoring" in offline analysis with python. This helped us in finding about a factor of sqrt(2) that we were missing earlier in 16171. we measured C1:ALS-BEATX_FINE_PHASE_OUT_HZ_DQ when X-ARM was locked to main laser and Xend green laser was locked to XARM. An excitation signal of amplitude 600 was setn at 619 hz at C1:ITMX_LSC_EXC.

Signal analysis flow:

  • The C1:ALS-BEATX_FINE_PHASE_OUT_HZ_DQ is calibrated to give value of beatntoe frequency in Hz. But we are interested in the fluctuations of this value at the excitation frequency. So the beatnote signal is first high passed with 50 hz cut-off. This value can be reduced a lot more in realtime system. We only took 60s of data and had to remove first 2 seconds for removing transients so we didn't reduce this cut-off further.
  • The I and Q demodulated beatntoe signal is combined to get a complex beatnote signal amplitude at excitation frequency.
  • This signal is divided by cts amplitude of excitation and multiplied by square of excitation frequency to get calibration factor for ITMX in units of nm/cts/Hz^2.
  • The noise spectrum of absolute value of  the calibration factor is plotted in attachment 1, along with its RMS. The calibration factor was detrended linearly so the the DC value was removed before taking the spectrum.
  • So Attachment 1 is the spectrum of noise in calibration factor when measured with this method. The shaded region is 15.865% - 84.135% percentile region around the solid median curves.

We got a value of \frac{7.3 \pm 3.9}{f^2}\, \frac{nm}{cts}.  The calibration factor in use is from \frac{7.32}{f^2} nm/cts from 13984.

Next steps could be to budget this noise while we setup some way of having this calibration factor generated in realitime using oscillators on a FE model. Calibrating actuation of a single optic in a single arm is easy, so this is a good test setup for getting a noise budget of this calibration method.

Attachment 1: ITMX_Cal_Noise_Spectrum_1307143423.pdf
ITMX_Cal_Noise_Spectrum_1307143423.pdf
  16191   Mon Jun 7 17:49:19 2021 Ian MacMillanUpdateCDSSUS simPlant model

Added difference to the graph. I included the code so that others could see what it looks like and use it for easy use.

Attachment 1: SingleSusPlantTF.pdf
SingleSusPlantTF.pdf
Attachment 2: TF_Graph_Code.zip
  16190   Mon Jun 7 15:37:01 2021 Anchal, Paco, YehonathanSummaryCamerasMon 7 in Control Room Died

We found Mon7 in control room dead today afternoon. It's front power on green light is not lighting up. All other monitors are working as normal.

This monitor was used for looking at IMC camera analog feed. It is one of the most important monitors for us, so we should replace it with a different monitor.

Yehonathan and Paco disconnected the monitor and brought it down. We put it under the back table if anyone wants to fix it. Paco has ordered a BNC to VGA/HDMI converter to put in any normal monitor up there. It will happen this Wednesday. Meanwhile, I have changed the MON4 assignment from POP to Quad2 to be used for IMC.

  16189   Mon Jun 7 13:14:20 2021 YehonathanUpdateCDSBI channels on c1auxey

I added a new XT1111 Acromag module to the c1auxey chassis. I sanitized and configured it according to the slow machines wiki instructions.

Since all the spare BIOs fit one DB37 connector I didn't add another feedthrough and combined them all on one and the same DB37 connector. This was possible because all the RTNs of the BIOs are tied to the chassis ground and therefore need only one connection. I changed the wiring spreadsheet accordingly.

I did a lot of rewirings and also cut short several long wires that were protruding from the chassis. I tested all the wires from the feedthroughs to the Acromag channels and fixed some wiring mistakes.

Tomorrow I will test the BIs using EPICs.

  16188   Sun Jun 6 16:33:47 2021 JonUpdateCDSBI channels on c1auxey

There is still an open issue with the BI channels not read by EPICS. They can still be read by the Windows machine though.

I looked into the issue that Yehonathan reported with the BI channels. I found the problem was with the .cmd file which sets up the Modbus interfacing of the Acromags to EPICS (/cvs/cds/caltech/target/c1auxey1/ETMYaux.cmd).

The problem is that all the channels on the XT1111 unit are being configured in Modbus as output channels. While it is possible to break up the address space of a single unit, so that some subset of channels are configured as inputs and another as outputs, I think this is likely to lead to mass confusion if the setup ever has to be modified. A simpler solution (and the convention we adopted for previous systems) is just to use separate Acromag units for BI and BO signals.

Accordingly, I updated the wiring plan to include the following changes:

  • The five EnableMon BI channels are moved to a new Acromag XT1111 unit (BIO01), whose channels are configured in Modbus as inputs.
  • One new DB37M connector is added for the 11 spare BI channels on BIO01.
  • The five channels freed up on the existing XT1111 (BIO00) are wired to the existing connector for spare BO channels.

So, one more Acromag XT1111 needs to be added to the c1auxey chassis, with the wiring changes as noted above. I have already updated the .cmd and EPICS database files in /cvs/cds/caltech/target/c1auxey1 to reflect these changes.

  16187   Sun Jun 6 15:59:51 2021 YehonathanUpdateCDSOpto-isolator for c1auxey

According to the BO interface circuit board https://dcc.ligo.org/D1001266, PCIN wires are connected to the coil driver and they are not pulled either way.

That means that they're either grounded or floating. I updated the drawing.

Quote:

This RTS also use the BO interface with an opto isolator. https://dcc.ligo.org/LIGO-D1002593

Could you also include the pull up/pull down situations?

 

Attachment 1: Optical_isolator_Wiring.pdf
Optical_isolator_Wiring.pdf
  16186   Sun Jun 6 12:15:16 2021 JonUpdateCDSOpto-isolator for c1auxey

Since this Ocean Controls optoisolator has been shown to be compatible, I've gone ahead and ordered 10 more:

  • (1) to complete c1auxey
  • (2) for the upgrade of c1auxex
  • (7) for the upgrade of c1susaux

They are expected to arrive by Wednesday.

  16185   Sun Jun 6 08:42:05 2021 JonUpdateCDSFront-End Assembly and Testing

Here is an update and status report on the new BHD front-ends (FEs).

Timing

The changes to the FE BIOS settings documented in [16167] do seem to have solved the timing issues. The RTS models ran for one week with no more timing failures. The IOP model on c1sus2 did die due to an unrelated "Channel hopping detected" error. This was traced back to a bug in the Simulink model, where two identical CDS parts were both mapped to ADC_0 instead of ADC_0/1. I made this correction and recompiled the model following the procedure in [15979].

Model naming standardization

For lack of a better name, I had originally set up the user model on c1sus2 as "c1sus2.mdl" This week I standardized the name to follow the three-letter subsystem convention, as four letters lead to some inconsistency in the naming of the auto-generated MEDM screens. I renamed the model c1sus2.mdl -> c1su2.mdl. The updated table of models is below.

Model Host CPU DCUID Path
c1x06 c1bhd 1 23 /opt/rtcds/userapps/release/cds/c1/models/c1x06.mdl
c1x07 c1sus2 1 24 /opt/rtcds/userapps/release/cds/c1/models/c1x07.mdl
c1bhd c1bhd 2 25 /opt/rtcds/userapps/release/isc/c1/models/c1bhd.mdl
c1su2 c1su2 2 26 /opt/rtcds/userapps/release/sus/c1/models/c1su2.mdl

Renaming an RTS model requires several steps to fully propagate the change, so I've documented the procedure below for future reference.

On the target FE, first stop the model to be renamed:

controls@c1sus2$ rtcds stop c1sus2

Then, navigate to the build directory and run the uninstall and cleanup scripts:

controls@c1sus2$ cd /opt/rtcds/caltech/c1/rtbuild/release
controls@c1sus2$ make uninstall-c1sus2
controls@c1sus2$ make clean-c1sus2

Unfortunately, the uninstall script does not remove every vestige of the old model, so some manual cleanup is required. First, open the file /opt/rtcds/caltech/c1/target/gds/param/testpoint.par and manually delete the three-line entry corresponding to the old model:

hostname=c1sus2
system=c1sus2
[C-node26]

If this is not removed, reinstallation of the renamed model will fail because its assigned DCUID will appear to already be in use. Next, find all relics of the old model using:

controls@c1sus2$ find /opt/rtcds/caltech/c1 -iname "*sus2*"

and manually delete each file and subdirectory containing the "sus2" name. Finally, rename, recompile, reinstall, and relaunch the model:

controls@c1sus2$ cd /opt/rtcds/userapps/release/sus/c1/models
controls@c1sus2$ mv c1sus2.mdl c1su2.mdl
controls@c1sus2$ cd /opt/rtcds/caltech/c1/rtbuild/release
controls@c1sus2$ make c1su2
controls@c1sus2$ make install-c1su2
controls@c1sus2$ rtcds start c1su2

Sitemap screens

I used a tool developed by Chris, mdl2adl, to auto-generate a set of temporary sitemap/model MEDM screens. This package parses each Simulink file and generates an MEDM screen whose background is an .svg image of the Simulink model. Each object in the image is overlaid with a clickable button linked to the auto-generated RTS screens. An example of the screen for the C1BHD model is shown in Attachment 1. Having these screens will make the testing much faster and less user-error prone.

I generated these screens following the instructions in Chris' README. However, I ran this script on the c1sim machine, where all the dependencies including Matlab 2021 are already set up. I simply copied the target .mdl files to the root level of the mdl2adl repo, ran the script (./mdl2adl.sh c1x06 c1x07 c1bhd c1su2), and then copied the output to /opt/rtcds/caltech/c1/medm/medm_teststand. Then I redefined the "sitemap" environment variable on the chiara clone to point to this new location, so that they can be launched in the teststand via the usual "sitemap" command.

Current status and plans

Is it possible to convert 18-bit AO channels to 16-bit?

Currently, we are missing five 18-bit DACs needed to complete the c1sus2 system (the c1bhd system is complete). Since the first shipment, we have had no luck getting additional 18-bit DACs from the sites, and I don't know when more will become available. So, this week I took an inventory of all the 16-bit DACs available at the 40m. I located four 16-bit DACs, pictured in Attachment 2. Their operational states are unknown, but none were labeled as known not to work.

The original CDS design would call for 40 more 18-bit DAC channels. Between the four 16-bit DACs there are 64 channels, so if only 3/4 of these DACs work we would have enough AO channels. However, my search turned up zero additional 16-bit DAC adapter boards. We could check if first Rolf or Todd have any spares. If not, I think it would be relatively cheap and fast to have four new adapters fabricated.

DAQ network limitations and plan

To get deeper into the signal-integrity aspect of the testing, it is going to be critical to get the secondary DAQ network running in the teststand. Of all the CDS tools (Ndscope, Diaggui, DataViewer, StripTool), only StripTool can be used without a functioning NDS server (which, in turn, requires a functioning DAQ server). StripTool connects directly to the EPICS server run by the RTS process. As such, StripTool is useful for basic DC tests of the fast channels, but it can only access the downsampled monitor channels. Ian and Anchal are going to carry out some simple DAC-to-ADC loopback tests to the furthest extent possible using StripTool (using DC signals) and will document their findings separately.

We don't yet have a working DAQ network because we are still missing one piece of critical hardware: a 10G switch compatible with the older Myricom network cards. In the older RCG version 3.x used by the 40m, the DAQ code is hardwired to interface with a Myricom 10G PCIe card. I was able to locate a spare Myricom card, pictured in Attachment 3, in the old fb machine. Since it looks like it is going to take some time to get an old 10G switch from the sites, I went ahead and ordered one this week. I have not been able to find documentation on our particular Myricom card, so it might be compatible with the latest 10G switches but I just don't know. So instead I bought exactly the same older (discontinued) model as is used in the 40m DAQ network, the Netgear GSM7352S. This way we'll also have a spare. The unit I bought is in "like-new" condition and will unfortunately take about a week to arrive.

Attachment 1: c1bhd.png
c1bhd.png
Attachment 2: 16bit_dacs.png
16bit_dacs.png
Attachment 3: myricom.png
myricom.png
  16184   Sun Jun 6 03:02:14 2021 KojiUpdateCDSOpto-isolator for c1auxey

This RTS also use the BO interface with an opto isolator. https://dcc.ligo.org/LIGO-D1002593

Could you also include the pull up/pull down situations?

  16183   Fri Jun 4 17:46:25 2021 unYehonathanUpdateCDSOpto-isolator for c1auxey

I mounted the optoisolator on the DIN rail and connected the 3 first channels

C1:SUS-ETMY_UL_ENABLE
C1:SUS-ETMY_UR_ENABLE

C1:SUS-ETMY_LL_ENABLE

to the optoisolator inputs 1,3,4 respectively. I connected the +15V input voltage into the input(+) of the optoisolator.

The outputs were connected to DB9F-2 where those channels were connected before.

I added DB9F-1 to the front panel to accept channels from the RTS. I connected the fast channels to connectors 1,2,3 from DB9F-1 to DB9F-2 according to the wiring diagram. The GND from DB9F-1 was connected to both connector 5 of DB9F-2 and the output (-).

I tested the channels: I connected a DB9 breakout board to DB9F-2. I measured the resistance between the RTS GND and the isolated channels while switching them on and off. In the beginning, when I turned on the binary channels the resistance was behaving weird - oscillating between low resistance and open circuit. I pulled up the channels through a 100Kohm resistor to observe whether the voltage behavior is reasonable or not. Indeed I observed that in the LOW state the voltage between the isolated channel and slow GND is 15V and 0.03V in the HIGH state. Then I disconnected the pull up from the channels and measured the resistance again. It showed ~ stable 170ohm in the HIGH state and an open circuit in the LOW state. I was not able to reproduce the weird initial behavior. Maybe the optoisolator needs some warmup of some sort.

 

We still need to wire the rest of the fast channels to DBF9-3 and isolate the channels in DBF9-4. For that, we need another optoisolator.

 

There is still an open issue with the BI channels not read by EPICS. They can still be read by the Windows machine though.

Attachment 1: 20210604_173420.jpg
20210604_173420.jpg
  16182   Fri Jun 4 14:49:23 2021 YehonathanUpdateCDSOpto-isolator for c1auxey

I made a diagram (Attached). I think it explains the blue thing in the previous post.

I don't know what is the grounding situation in the RTS so I put a ground in both the coil driver and the RTS. Hopefully, only one of them is connected in reality.

Quote:

- Could you explain what is the blue thing in Attachment 1?

- To check the validity of the signal chain, can you make a diagram summarizing the path from the fast BO - BO I/F - Acromag - This opto-isolator - the coil driver relay? (Cut-and-paste of the existing schematics is fine)

 

 

Attachment 1: Optical_isolator_Wiring.pdf
Optical_isolator_Wiring.pdf
  16181   Thu Jun 3 22:08:00 2021 KojiUpdateCDSOpto-isolator for c1auxey

- Could you explain what is the blue thing in Attachment 1?

- To check the validity of the signal chain, can you make a diagram summarizing the path from the fast BO - BO I/F - Acromag - This opto-isolator - the coil driver relay? (Cut-and-paste of the existing schematics is fine)

 

  16180   Thu Jun 3 17:49:46 2021 PacoUpdateEquipment loanBorrow red cart

Returned today.

Quote:

I borrowed the little red cart 🛒 to help clear the path for new optical tables in B252 West Bridge. Will return once I am done with it.  

 

  16179   Thu Jun 3 17:35:31 2021 AnchalSummaryIMCFixed medm button

I fixed the PSL shutter button on Shutters summary page C1IOO_Mech_Shutter.adl. Now PSL switch changes C1:PSL-PSL_ShutterRqst channel. Earlier it was C1:AUX-PSL_ShutterRqst which doesn't do anything.

 

Attachment 1: C1IOO_Mech_Shutters.png
C1IOO_Mech_Shutters.png
  16178   Thu Jun 3 17:15:17 2021 YehonathanUpdateCDSOpto-isolator for c1auxey

As Jon wrote we need to use the NPN configuration (see attachments). I tested the isolator channels in the following way:

1. I connected +15V from the power supply to the input(+) contact.

2. Signal wire from one of the digital outputs was connected to I1-4

3. When I set the digital output to HIGH, the LED on the isolator turns on.

4. I measure the resistance between O1-4 to output(-) and find it to be ~ 100ohm in the HIGH state and an open circuit in the LOW state, as expected from an open collector output.

Unlike the Acromag output, the isolator output is not pulled up in the LOW state. To do so we need to connect +15V to the output channel through a pull-up resistor. For now, I leave it with no pull-up. According to the schematics of the HAM-A Coil Driver, the digital output channels drive an electromagnetic relay (I think) so it might not need to be pulled up to switch back. I'm not sure. We will need to check the operation of these outputs at the installation.

During the testing of the isolator outputs pull-up, I accidentally ran a high current through O2, frying it dead. It is now permanently shorted to the + and - outputs rendering it unusable. In any case, we need another isolator since we have 5 channels we need to isolate.

I mounted the isolator on the DIN rail and started wiring the digital outputs into it. I connected the GND from the RTS to output(-) such that when the digital outputs are HIGH the channels in the coil driver will be sunk into the RTS GND and not the slow one avoiding GND contamination.

Attachment 1: Optical_Isolator_NPN_Input.png
Optical_Isolator_NPN_Input.png
Attachment 2: Optical_Isolator_NPN_Output.png
Optical_Isolator_NPN_Output.png
  16177   Thu Jun 3 13:06:47 2021 Ian MacMillanUpdateCDSSUS simPlant model

I was able to measure the transfer function of the plant filter module from the channel X1:SUP-C1_SUS_SINGLE_PLANT_Plant_POS_Mod_EXC to X1:SUP-C1_SUS_SINGLE_PLANT_Plant_POS_Mod_OUT. The resulting transfer function is shown below. I have also attached the raw data for making the graph.

Next, I will make a script that will make the photon filters for all the degrees of freedom and start working on the matrix version of the filter module so that there can be multiple degrees of freedom.

Attachment 1: SingleSusPlantTF.pdf
SingleSusPlantTF.pdf
Attachment 2: SUS_PLANT_TF.zip
  16176   Wed Jun 2 17:50:50 2021 PacoUpdateEquipment loanBorrow red cart

I borrowed the little red cart 🛒 to help clear the path for new optical tables in B252 West Bridge. Will return once I am done with it.  

Attachment 1: IMG_20210602_172858.jpg
IMG_20210602_172858.jpg
  16175   Wed Jun 2 16:20:59 2021 Anchal, PacoSummarySUSIMC Suspension gains reverted to old values

Following the conclusion, we are reverting the suspension gains to old values, i.e.

IMC Suspension Gains
  MC1 MC2 MC3
SUSPOS 120 150 200
SUSPIT 60 10 12
SUSYAW 60 10 8

While the F2A filters, AC coil gains and input matrices are changed to as mentioned in 16066 and 16072.

The changes can be reverted all the way back to old settings (before Paco and I changed anything in the IMC suspensions) by running python scripts/SUS/general/20210602_NewIMCOldGains/restoreOldConfigIMC.py on allegra. The new settings can be uploaded back by running python scripts/SUS/general/20210602_NewIMCOldGains/uploadNewConfigIMC.py on allegra.


Change time:

Unix Time = 1622676038

UTC Jun 02, 2021 23:20:38 UTC
Central Jun 02, 2021 18:20:38 CDT
Pacific Jun 02, 2021 16:20:38 PDT

GPS Time = 1306711256

Quote:
 
  • Conclusion 1:  The 'new settings with new gains' cause more coupling to seismic noise, probably due to low phase margin in control loops. We should revert back the suspension damping gains.
  • Conclusion 2: The 'new settings' work as expected and can be kept when WFS loops are optimized further.
  • Conjecture: From our experience over last 2 weeks, locking the arms to the main laser with 'new settings with new gains' introduces noise in the arm length large enough that the Xend green laser does not remain locked to the arm for longer than tens of seconds. So this is definitely not a configuration in which we can carry out other measurements and experiments in the interferometer.

 

  16174   Wed Jun 2 09:43:30 2021 Anchal, PacoSummarySUSIMC Settings characterization

Plot description:

  • We picked up three 10 min times belonging to the three different configurations:
    • 'Old Settings': IMC Suspension settings before Paco and I changed anything. Data taken from Apr 26, 2021, 00:30:42 PDT (GPS 1303457460).
    • 'New Settings': New input matrices uploaded on April 28th, along with F2A filters and AC coil balancing gains (see 16091). Data taken from May 01, 2021, 00:30:42 PDT (GPS 1303889460).
    • 'New settings with new gains' Above and new suspension damping gains uploaded on May5th, 2021 (see 16120). Data taken from May 07, 2021, 03:10:42 PDT (GPS 1304417460).
  • Attachment 1  shows the RMS seismic noise along X direction between 1 Hz and 3 Hz picked from C1:PEM-RMS_BS_X_1_3 during the three time durations chosen. This plot is to establish that RMS noise levels were similar and mostly constant. Page 2 shows the mean ampltidue spectral density of seismic noise in x-direction over the 3 durations.
  • Attachment 2 shows the transfer function estimate of seismic noise to MC_F during the three durations. Page 1 shows ratio of ASDs taken with median averaging while page 2 shows the same for mean averaging.
  • Attachment 3 shows the transfer function estimate of seismic noise to MC_TRANS_PIT during the three durations. Page 1 shows ratio of ASDs taken with median averaging while page 2 shows the same for mean averaging.
  • Attachment 4 shows the transfer function estimate of seismic noise to MC_TRANS_YAW during the three durations. Page 1 shows ratio of ASDs taken with median averaging while page 2 shows the same for mean averaging.

Inferences:

  • From Attachment 2 Page 1:
    • We see that 'old settings' caused the least coupling of seismic noise to MC_F signal in most of the low frequency band except between 1.5 to 3 Hz where the 'new settings' were slightly better.
    • 'new settings' also show less coupling in 4 Hz to 6 Hz band, but at these frequencies, seismix noise is filtered out by suspension, so this could be just coincidental and is not really a sign of better configuration.
    • There is excess noise coupling seen with 'new settings' between 0.4 Hz and 1.5 Hz. We're not sure why this coupling increased.
    • 'new settings with new gains' show the most coupling in most of the frequency band. Clearly, the increased suspension damping gains did not behaved well with rest of the system.
  • From Attachment 3 Page 1:
    • Coupling to MC_TRANS_PIT error signal is reduced for 'new settings' in almost all of the frequency band in comparison to the 'old settings'.
    • 'new settings with new gains' did even better below 1 Hz but had excess noise in 1 Hz to 6 Hz band. Again increased suspension damping gains did not help much.
    • But low coupling to PIT error for 'new settings' suggest that our decoupling efforts in matrix diagonalization, F2A filters and ac coil balancing worked to some extent.
  • From Attachment 4 Page 1:
    • 'new settings' and 'old settings' have the same coupling of seismic noise to MC_TRANS_YAW in all of the frequency band. This is in-line witht eh fact that we found very little POS to YAW couping in our analysis before and there was little to no change for these settings.
    • 'new settings with new gains' did better below 1 Hz but here too there was excess coupling between 1 Hz to 9 Hz.
  • Page 1 vs Page 2:
    • Mean and median should be same if the data sample was large enough and noise was stationary. A difference between the two suggests existence of outliers in the data set and median provides a better central estimate in such case.
    • MC_F: Mean and median are same below 4 hz. There are high frequency outliers above 4 Hz in 'new settings with new gains' and 'old settings' data sets, maybe due to transient higher free running laser frequency noise. But since, suspension settigns affect below 1 Hz mostly, the data sets chosen are stationary enough for us.
    • MC_TRANS_PIT: Mean ratio is lower for 'new settings' and 'old settings' in 0.3 hz to 0.8 Hz band. Same case above 4 Hz as listed above.
    • MC_TRANS_YAW:  Same as above.
  • Conclusion 1:  The 'new settings with new gains' cause more coupling to seismic noise, probably due to low phase margin in control loops. We should revert back the suspension damping gains.
  • Conclusion 2: The 'new settings' work as expected and can be kept when WFS loops are optimized further.
  • Conjecture: From our experience over last 2 weeks, locking the arms to the main laser with 'new settings with new gains' introduces noise in the arm length large enough that the Xend green laser does not remain locked to the arm for longer than tens of seconds. So this is definitely not a configuration in which we can carry out other measurements and experiments in the interferometer.
Attachment 1: seismicX.pdf
seismicX.pdf seismicX.pdf
Attachment 2: seismicXtoMC_F_TFest.pdf
seismicXtoMC_F_TFest.pdf seismicXtoMC_F_TFest.pdf
Attachment 3: seismicXtoMC_TRANS_PIT_TFest.pdf
seismicXtoMC_TRANS_PIT_TFest.pdf seismicXtoMC_TRANS_PIT_TFest.pdf
Attachment 4: seismicXtoMC_TRANS_YAW_TFest.pdf
seismicXtoMC_TRANS_YAW_TFest.pdf seismicXtoMC_TRANS_YAW_TFest.pdf
  16173   Wed Jun 2 01:08:57 2021 KojiUpdateSUSCoM to Clamping Point Measurement for 3" Adapter Ring

How about to use the non-Ag coated threaded shaft + the end SS masses with helicoils inserted? Does this save the masses to get stuck?

 

  16172   Wed Jun 2 01:03:19 2021 KojiUpdateBHDSOS assembly

Can you just cut the viton tips smaller? If you cut it to have some wedge (or say, taper), it can get stuck with the vent hole.

 

  16171   Tue Jun 1 16:55:32 2021 Anchal, PacoSummaryALSSingle Arm Actuation Calibration with IR ALS Beat

Rana suggested in today's meeting to put in a notch filter in the XARM IR PDH loop to avoid suppressing the excitation line. We tried this today first with just one notch at 1069 Hz and then with an additional notch at 619 Hz and sent two simultaneous excitations.


Measurement and Analysis:

  • We added notch filters with Q=10, depth=50dB, freq=619 Hz and 1069 Hz using foton in SUS-ETMX_LSC filter bank at FM10.
  • We sent excitation signals with amplitudes 600cts and 1000 cts for 619 Hz and 1069 Hz signals respectively.
  • We measured time series data of C1:SUS-ITMX_LSC_OUT_DQ and C1:ALS-BEATX_FINE_PHASE_OUT_HZ_DQ for 60s.
  • Then, spectrum of both signals is measured with Hanning window using scipy.welch function with scaling set to  'spectrum', binwidth=1Hz.
  • The beatnote signal was converted into length units by multiplying it by 1064nm * 37.79m / c.
  • The ratio of the two spectrums at teh excitation frequency multiplies by excitation frequency squared gives us teh calibration constant in units of nm Hz^2/cts.
  • At 619 Hz, we got \frac{5.01}{f^2}nm/cts
  • At 1069 Hz, we got \frac{5.64}{f^2}nm/cts.
  • The calibration factor in use is from \frac{7.32}{f^2} nm/cts from 13984.
  • So, the calibration factor from this methos is about 23% smaller than measured using freeswinging MICH in 13984.
  • One possiblity is that our notch filter is not as effective in avoiding suppresion of excitation.
    • We tried increasing the notch filter depths to 100 dB but got the same result within 2%.
    • We tried changing the position of notch filters. We put them in POX filter banks. Again the result did not change more than 2%.
  • The open loop gain of green PDH at 619 Hz and 1069 Hz must be large enough for our assumption of green laser perfectly following length motion to be true. The UGF of green laser is near 11 kHz.
  • The discrepancy could be due to outdated freeswinging MICH measurement that was done 3 years ago. Maybe we should learn how to do the ITMX calibration using this method and compare our own two measurements.
Attachment 1: SingleArmActCalwithIRALSBeat-1306624785.pdf
SingleArmActCalwithIRALSBeat-1306624785.pdf
  16170   Tue Jun 1 16:17:06 2021 YehonathanUpdateBHDSOS assembly

I tried to push the clean Viton tips into the vented screws just to find out that the vented holes are too small. We need to drill 0.1" diameter holes about 0.1" deep into these screws and clean them again.

 

  16169   Tue Jun 1 14:26:23 2021 JordanUpdateSUSCoM to Clamping Point Measurement for 3" Adapter Ring

After changing the material of the Balance Mass from 6061 Al to 304 Steel, and changing the thickness to 0.21" from 0.25". The CoM is now 1.11mm below the clamping point.

Koji expected a mass change of ~ 4g to move the mass to 1.1mm. The 6061 mass weighed ~1.31g and the 304 mass weighs 4.1g.

A potential issue with this is the screw used the adjust the position of these balance masses, threads through both the aluminum ring and this now 304 steel mass. A non silver plated screw could cold weld at the mass, but a silver plated screw will gall in the aluminum threads.

Quote:

The current vertical distance between the CoM and the wire clamping point on the 3" Ring assembly is 0.33mm. That is the CoM is .33 mm below the clamping point of the wire. I took the clamping point to be the top edge of the wire clamp piece. see the below attachments.

I am now modifying the dumbell mechanism at the bottom of the ring to move the CoM to the target distance of 1.1mm.

 

Attachment 1: CoM_to_Clamp_Updated.PNG
CoM_to_Clamp_Updated.PNG
  16168   Fri May 28 17:32:48 2021 AnchalSummaryALSSingle Arm Actuation Calibration with IR ALS Beat

I attempted a single arm actuation calibration using IR beatnote (in the directions of soCal idea for DARM calibration)


Measurement and Inferences:

  • I sent 4 excitation signals at C1:SUS-ITM_LSC_EXC wit 30cts at 31Hz, 200cts at 197Hz, 600cts at 619Hz and 1000cts at 1069 Hz.
  • These were sent simultaneously using compose function in python awg.
  • The XARM was locked to mai laser and alignment was optimized with ASS.
  • The Xend Green laser was locked to XARM and alignment was optimized.
    • Sidenote: GTRX is now normalized to give 1 at near maximum power.
    • Green lasers can be locked with script instead of toggling.
    • Script can be called from sitemap->ALS->! Toggle Shutters->Lock X Green
    • Script is present at scripts/ALS/lockGreen.py.
  • C1:ALS-BEATX_FINE_PHASE_OUT_HZ_DQ was measured for 60s.
  • Also, measured C1:LSC-XARM_OUT_DQ and C1:SUS-ITMX_LSC_OUT_DQ.
  • Attachment 1 shows the measured beatnote spectrum with excitations on in units of m/rtHz.
  • It also shows resdiual displacement contribution PSD of (output referred) XARM_OUT and ITMX_LSC_OUT to the same point in the state space model.
    • Note: that XARM_OUT and ITMX_LSC_OUT (excitation signal) get coherently added in reality and hence the beatnote spectrum at each excitation frequency is lower than both of them.
    • The remaining task is to figure out how to calculate the calibration constant for ITMX actuation from this information.
    • I need more time to understand the mixture of XARM_OUT and ITMX_LSC_OUT in the XARM length node in control loop.
    • Beatnote signal tells us the actual motion of the arm length, not how much ITMX would have actuated if the arm was not locked.
  • Attachment 2 has the A,B,C,D matrices for the full state space model used. These were fed to python controls package to get transfer functions from one point to another in this MIMO.
    • Note, that here I used the calibration of XARM_OUT we measured earlies in 16127.
    • On second thought, maybe I should first send excitation in ETMX_LSC_EXC. Then, I can just measure ETMX_LSC_OUT which includes XARM_OUT due to the lock and use that to get calibration of ETMX actuation directly.

Attachment 1: SingleArmActCalwithIRALSBeat.pdf
SingleArmActCalwithIRALSBeat.pdf
Attachment 2: stateSpaceModel.zip
  16167   Fri May 28 11:16:21 2021 JonUpdateCDSFront-End Assembly and Testing

An update on recent progress in the lab towards building and testing the new FEs.

1. Timing problems resolved / FE BIOS changes

The previously reported problem with the IOPs losing sync after a few minutes (16130) was resolved through a change in BIOS settings. However, there are many required settings and it is not trivial to get these right, so I document the procedure here for future reference.

The CDS group has a document (T1300430) listing the correct settings for each type of motherboard used in aLIGO. All of the machines received from LLO contain the oldest motherboards: the Supermicro X8DTU. Quoting from the document, the BIOS must be configured to enforce the following:

• Remove hyper-threading so the CPU doesn’t try to run stuff on the idle core, as hyperthreading simulate two cores for every physical core.
• Minimize any system interrupts from hardware, such as USB and Serial Ports, that might get through to the ‘idled’ core. This is needed on the older machines.
• Prevent the computer from reducing the clock speed on any cores to ‘save power’, etc. We need to have a constant clock speed on every ‘idled’ CPU core.

I generally followed the T1300430 instructions but found a few adjustments were necessary for diskless and deterministic operation, as noted below. The procedure for configuring the FE BIOS is as follows:

  1. At boot-up, hit the delete key to enter the BIOS setup screen.
  2. Before changing anything, I recommend photographing or otherwise documenting the current working settings on all the subscreens, in case for some reason it is necessary to revert.
  3. T1300430 assumes the process is started from a known state and lists only the non-default settings that must be changed. To put the BIOS into this known state, first navigate to Exit > Load Failsafe Defaults > Enter.
  4. Configure the non-default settings following T1300430 (Sec. 5 for the X8DTU motherboard). On the IPMI screen, set the static IP address and netmask to their specific assigned values, but do set the gateway address to all zeros as the document indicates. This is to prevent the IPMI from trying to initiate outgoing connections.
  5. For diskless booting to continue to work, it is also necessary to set Advanced > PCI/PnP Configuration > Load Onboard LAN 1 Option Rom > Enabled.
  6. I also found it was necessary to re-enable IDE direct memory access and WHEA (Windows Hardware Error Architecture) support. Since these machines have neither hard disks nor Windows, I have no idea why these are needed, but I found that without them, one of the FEs would hang during boot about 50% of the time.
    • Advanced > PCI/PnP configuration > PCI IDE BusMaster  > Enabled.
    • Advanced > ACPI Configuration > WHEA Support > Enabled.

After completing the BIOS setup, I rebooted the new FEs about six times each to make sure the configuration was stable (i.e., would never hang during boot).

2. User models created for FE testing

With the timing issue resolved, I proceeded to build basic user models for c1bhd and c1sus2 for testing purposes. Each one has a simple structure where M ADC inputs are routed through IIR filters to an output matrix, which forms linear signal combinations that are routed to N DAC outputs. This is shown in Attachment 1 for the c1bhd case, where the signals from a single ADC are conditioned and routed to a single 18-bit DAC. The c1sus2 case is similar; however the Contec BO modules still needed to be added to this model.

The FEs are now running two models each: the IOP model and one user model. The assigned parameters of each model are documented below.

Model Host CPU DCUID Path
c1x06 c1bhd 1 23 /opt/rtcds/userapps/release/cds/c1/models/c1x06.mdl
c1x07 c1sus2 1 24 /opt/rtcds/userapps/release/cds/c1/models/c1x07.mdl
c1bhd c1bhd 2 25 /opt/rtcds/userapps/release/isc/c1/models/c1bhd.mdl
c1sus2 c1sus2 2 26 /opt/rtcds/userapps/release/sus/c1/models/c1sus2.mdl

The user models were compiled and installed following the previously documented procedure (15979). As shown in Attachment 2, all the RTS processes are now working, with the exception of the DAQ server (for which we're still awaiting hardware). Note that these models currently exist only on the cloned copy of the /opt/rtcds disk running on the test stand. The plan is to copy these models to the main 40m disk later, once the new FEs are ready to be installed.

3. AA and AI chassis installed

I installed several new AA and AI chassis in the test stand to interface with the ADC and DAC cards. This includes three 16-bit AA chassis, one 16-bit AI chassis, and one 18-bit AI chassis, as pictured in Attachment 3. All of the AA/AI chassis are powered by one of the new 15V DC power strips connected to a bench supply, which is housed underneath the computers as pictured in Attachment 4.

These chassis have not yet been tested, beyond verifying that the LEDs all illuminate to indicate that power is present.

Attachment 1: c1bhd.png
c1bhd.png
Attachment 2: gds_tp.png
gds_tp.png
Attachment 3: teststand.jpeg
teststand.jpeg
Attachment 4: bench_supply.jpeg
bench_supply.jpeg
  16166   Fri May 28 10:54:59 2021 JonUpdateCDSOpto-isolator for c1auxey

I have received the opto-isolator needed to complete the new c1auxey system. I left it sitting on the electronics bench next to the Acromag chassis.

Here is the manufacturer's wiring manual. It should be wired to the +15V chassis power and to the common return from the coil driver, following the instructions herein for NPN-style signals. Note that there are two sets of DIP switches (one on the input side and one on the output side) for selecting the mode of operation. These should all be set to "NPN" mode.

Attachment 1: optoisolator.jpeg
optoisolator.jpeg
  16165   Thu May 27 14:11:15 2021 JordanUpdateSUSCoM to Clamping Point Measurement for 3" Adapter Ring

The current vertical distance between the CoM and the wire clamping point on the 3" Ring assembly is 0.33mm. That is the CoM is .33 mm below the clamping point of the wire. I took the clamping point to be the top edge of the wire clamp piece. see the below attachments.

I am now modifying the dumbell mechanism at the bottom of the ring to move the CoM to the target distance of 1.1mm.

Attachment 1: CoM_to_Clamp.PNG
CoM_to_Clamp.PNG
Attachment 2: CoM_to_Clamp_2.PNG
CoM_to_Clamp_2.PNG
  16164   Thu May 27 11:03:15 2021 Anchal, PacoSummaryALSALS Single Arm Noise Budget

Here's an updated X ARM ALS noise budget.

Things to remember:

  • Major mistake we were making earlier was that we were missing the step of clicking  'Set Phase UGF' before taking the measurement.
  • Click the clear phase history just before taking measure.
  • Make sure the IR beatnotes are less than 50 MHz (or the left half of HP8591E on water crate). The DFD is designed for this much beatnote frequency (from Gautum).
  • We took this measurement with old IMC settings.
  • We have saved a template file in users/Templates/ALS/ALS_outOfLoop_Ref_DQ.xml . This si same as ALS_outOfLoop_Ref.xml except we changed all channels to _DQ.

Conclusions:

  • Attachment 1 shows the updated noisebudget. The estimated and measured RMS noise are very close to eachother.
  • However, there is significant excess noise between 4 Hz and 200 Hz. We're still thinking on what could be the source of these.
  • From 200 Hz to about 3 kHz, the beatnote noise is dominated by AUX residual frequency noise. This can be verified with page 2 of Attachment 2 where coherence between AUX PDH Error signal and BEATX signal is high.
  • One mystery is how the measured beatnote noise is below the residual green laser noise above 3 kHz. Could this be just because the phase tracker can't measure noise above 3kHz?
  • We have used estimated open loop transfer function for AUX from poles/zeros for uPDH box used (this was done months ago by me when I was working on ALS noise budget from home). We should verify it with a fresh OLTF measurement of AUX PDH loop. That's next on our list.
Attachment 1: ALS_Single_X_Arm_IR.pdf
ALS_Single_X_Arm_IR.pdf
Attachment 2: ALS_OOL_with_Ref.pdf
ALS_OOL_with_Ref.pdf ALS_OOL_with_Ref.pdf ALS_OOL_with_Ref.pdf ALS_OOL_with_Ref.pdf
  16163   Wed May 26 11:45:57 2021 Anchal, PacoConfigurationIMCMC2 analog camera

[Anchal, Paco]

We went near the MC2 area and opened the lid to inspect the GigE and analog video monitors for MC2. Looked like whatever image is coming through the viewport is split into the GigE (for beam tracking) and the analog monitor. We hooked the monitor found on the floor nearby and tweaked the analog video camera around to get a feel for how the "ghost" image of the transmission moves around. It looks like in order to try and remove this "extra spots" we would need to tweak the beam tracking BS. We will consult the beam tracking authorities and return to this.

  16162   Wed May 26 02:00:44 2021 gautamUpdateElectronicsCoil driver noise

I was preparing a short write-up / test procedure for the custom HV coil driver, when I thought of something I can't resolve. I'm probably missing some really basic physics here - but why do we not account for the shot noise from DC current flowing through the series resistor? For a 4kohm resistor, the Johnson current noise is ~2pA/rtHz. This is the target we were trying to beat with our custom designed HV bias circuit. But if there is a 1 mA DC current flowing through this resistor, the shot noise of this current is \sqrt{2eI_{\mathrm{DC}}} \approx18pA/rtHz, which is ~9 times larger than the Johnson noise of the same resistor. One could question the applicability of this formula to calculate the shot noise of a DC current through a wire-wound resistor - e.g. maybe the electron transport is not really "ballistic", and so the assumption that the electrons transported through it are independent and non-interacting isn't valid. There are some modified formulae for the shot noise through a metal resistor, which evaluates to \sqrt{2eI_{\mathrm{DC}}/3} \approx10pA/rtHz for the same 4kohm resistor, which is still ~5x the Johnson noise. 

In the case of the HV coil driver circuit, the passive filtering stage I added at the output to filter out the excess PA95 noise unwittingly helps us - the pole at ~0.7 Hz filters the shot noise (but not the Johnson noise) such that at ~10 Hz, the Johnson noise does indeed dominate the total contribution. So, for this circuit, I think we don't have to worry about some un-budgeted noise. However, I am concerned about the fast actuation path - we were all along assuming that this path would be dominated by the Johnson noise of the 4kohm series resistor. But if we need even 1mA of current to null some DC DARM drift, then we'd have the shot noise contribution become comparable, or even dominant?

I looked through the iLIGO literature, where single-stage suspensions were being used, e.g. Rana's manifesto, but I cannot find any mention of shot noise due to DC current, so probably there is a simple explanation why - but it eludes me, at least for the moment. The iLIGO coil drivers did not have a passive filter at the output of the coil driver circuit (at least, not till this work), and there isn't any feedback gain for the DARM loop at >100 Hz (where we hope to measure squeezing) to significantly squash this noise.

Attachment #1 shows schematic topologies of the iLIGO and proposed 40m configs. It may be that I have completely misunderstood the iLIGO config and what I've drawn there is wrong. Since we are mainly interested in the noise from the resistor, I've assumed everything upstream of the final op-amp is noiseless (equivalently, we assume we can sufficiently pre-filter these noises).
Attachment #2 shows the relative magnitudes of shot noise due to a DC current, and thermal noise of the series resistor, as a function of frequency, for a few representative currents, for the slow bias path assuming a 0.7Hz corner from the 4kohm/3uF RC filter at the output of the PA95.


Some lit review suggests that it's actually pretty hard to measure shot noise in a resistor - so I'm guessing that's what it is, the mean free path of electrons is short compared to the length of the resistor such that the assumption that electrons arrive independently and randomly isn't valid. So Ohm's law dictates I=V/R and that's what sets the current noise. See, for example, pg 432 of Horowitz and Hill.

Attachment 1: coilDriverTopologies.pdf
coilDriverTopologies.pdf
Attachment 2: shotVthermal.pdf
shotVthermal.pdf
  16161   Tue May 25 17:42:11 2021 Anchal, PacoSummaryALSALS Single Arm Noise Budget

Here is our first attempt at a single-arm noise budget for ALS.

Attachment 1 shows the loop diagram we used to calculate the contribution of different noises.

Attachment 2 shows the measured noise at C1:ALS-BEATX_PHASE_FINE_OUT_HZ when XARM was locked to the main laser and Xend Green laser was locked to XARM.

  • The brown curve shows the measured noise.
  • The black curve shows total estimated noise from various noise sources (some of these sources have not been plotted as their contribution falls off the plotting y-lims.)
  • The residual frequency noise of Xend green laser (AUX) is measured by measuring the PDH error monitor spectrum from C1:ALS-X_ERR_MON_OUT_DQ. This measurement was converted into units of V by multiplying it by 6.285e-4 V/cts. This factor was measured by sending a 43 Hz 100 mV sine wave at the readout point and measuring the output in the channel.
  • This error signal is referred to AUX_Freq input in the loop diagram (see attachment 1) and then injected from there.
  • All measurements were taken to Res_Disp port in the 'Out-of-Loop Beat Note' block (see attachment 1).
  • In this measurement, we did not DAC noise that gets added when ALS loop is closed.
  • We added ADC noise from Kiwamu's ALS paper after referring it to DFD input. DFD noise is also taken from Kiwamu's ALS paper data.

Inference:

  • Something is wrong above 200 Hz for the inclusion of AUX residual displacement noise. It is coming off as higher than the direct measured residual noise, so something is wrong with our loop diagram. But I'm not sure what.
  • There is a lot of unaccounted noise everywhere from 1 Hz to 200 Hz.
  • Rana said noise budget below 1 Hz is level 9 stuff while we are at level 2, so I'll just assume the excess noise below 1 Hz is level 9 stuff.
  • We did include seismic noise taken from 40m noise budget in 40m/pygwinc. But it seems to affect below the plotted ylims. I'm not sure if that is correct either.

Unrelated questions:

  • There is a slow servo feeding back to Green Laser's crystal temperature by integrating PZT out signal. This is OFF right now. Should we keep it on?
  • The green laser lock is very unreliable and it unlocks soon after any signal is being fed back to the ETMX position.
  • This means, keeping both IR and green light locked in XARM is hard and simultaneous oscillation does not last longer than 10s of seconds. Why is it like this?
  • We notice that multiple higher-order modes from the green laser reach the arm cavity. The HOMs are powerful enough that PDH locks to them as well and we toggle the shutter to come to TEM00 mode. These HOMs must be degrading the PDH error signal. Should we consider installing PMCs at end tables too?
Attachment 1: ALS_IR_b.svg
ALS_IR_b.svg
Attachment 2: ALS_Single_Arm_IR.pdf
ALS_Single_Arm_IR.pdf
  16160   Tue May 25 17:08:17 2021 ChubUpdateElectronicschassis rework complete!

All remaining chasses have been reworked and placed on the floor along the west wall in Room 104. 

Attachment 1: 40M_chassis_reworked_5-25-21.jpg
40M_chassis_reworked_5-25-21.jpg
  16159   Tue May 25 10:22:16 2021 Anchal, PacoSummarySUSMC1 new input matrix calculated and uploaded

The test was succesful and brought back the IMC to lock point at the end.

We calculated new input matrix using same code in scripts/SUS/InMatCalc/sus_diagonalization.py . Attachment 1 shows the results.

The calculations are present in scripts/SUS/InMatCalc/MC1.


We uploaded the new MC1 input matrix at:

Unix Time = 1621963200

UTC May 25, 2021 17:20:00 UTC
Central May 25, 2021 12:20:00 CDT
Pacific May 25, 2021 10:20:00 PDT

GPS Time = 1305998418

This was done by running python scripts/SUS/general/20210525_NewMC1Settings/uploadNewConfigIMC.py on allegra. Old IMC settings (before Paco and I started workin on 40m) can be restored by running python scripts/SUS/general/20210525_NewMC1Settings/restoreOldConfigIMC.py on allegra.

Everything looks as stable as before. We'll look into long term trends in a week to see if this helped at all.

Attachment 1: SUS_Input_Matrix_Diagonalization.pdf
SUS_Input_Matrix_Diagonalization.pdf
  16158   Mon May 24 20:55:00 2021 KojiSummaryBHDHow to align two OMCs on the BHD platform?

Differential misalignment of the OMCs

40m BHD will employ two OMCs on the BHD platform. We will have two SOSs for each of the LO and AS beams. The challenge here is that the input beam must optimally couple to the OMCs simultaneously. This is not easy as we won't have independent actuators for each OMC. e.g. The alignment of the LO beam can be optimally adjusted to the OMC1, but this, in general, does not mean the beam is optimally aligned to the OMC2.

Requirement

When a beam with the matched mode to an optical cavity has a misalignment, the power coupling C can be reduced from the unity as

C = 1 - \left(\frac{a}{\omega_0}\right)^2 - \left(\frac{\alpha}{\theta_0}\right)^2

where \omega_0 is the waist radius, \theta_0 is the divergence angle defined as \theta_0 \equiv \lambda/ \pi \omega, a and \alpha are the beam lateral translation and rotation at the waist position.

The waist size of the OMC is 500um. Therefore \omega_0 = 500um and \theta_0 = 0.68 mrad. If we require C to be better than 0.995 according to the design requirement document (T1900761). This corresponds to a (only) to be 35um and \alpha (only) to be 48urad. These numbers are quite tough to be realized without post-installation adjustment. Moreover, the OMCs themselves have individual differences in the beam axis. So no matter how we set the mechanical precision of the OMC installation, we will introduce a maximum of 1mm and ~5mrad uncertainty of the optical axis.

Adjustment

Suppose we adjust the incident beam to the OMC placed at the transmission side of the BHD BS. The reflected beam at the BS can be steered by picomotors. The distance from the BS to the OMC waist is 12.7" (322mm) according to the drawing.
So we can absorb the misalignment mode of (a, \alpha) = (0.322 \theta, \theta). This is a bit unfortunate. 0.322m is about 1/2 of the rayleigh range. Therefore, this actuation is still angle-dominated but a bit of translation is still coupled.

If we enable to use the third picomotor on the BHD BS mount, we can introduce the translation of the beam in the horiz direction. This is not too huge therefore we still want to prepare the method to align the OMC in the horiz direction.

The difficult problem is the vertical alignment. This requires the vertical displacement of the OMC. And we will not have the option to lower the OMC. Therefore if the OMC2 is too high, we have to raise the OMC1 so that the resulting beam is aligned to the OMC2. i.e. we need to maintain the method to raise both OMCs. (... or swap the OMCs). From the images of the OMC beam spots, we'll probably be able to analyze the intracavity axes of the OMCs. So we can always place the OMC with a higher optical axis at the transmission side of the BHD BS.

 

 

  16157   Mon May 24 19:14:15 2021 Anchal, PacoSummarySUSMC1 Free Swing Test set to trigger

We've set a free swing test to trigger at 3:30 am tomorrow for MC1. The script for tests is running on tmux session named 'freeSwingMC1' on rossa. The script will run for about 4.5 hrs and we'll correct the input matrix tomorrow from the results. If anyone wants to work during this time (3:30 am to 8:00 am), you can just kill the script by killing tmux session on rossa. ssh into rossa and type tmux kill-session -t freeSwingMC1.

Quote:
 

We should redo the MC1 input matrix optimization and the coil balancing afterward as we did everything based on the noisy UL OSEM values.

 

ELOG V3.1.3-